marcia is a fashion designer who runs a small business in a competitive industry. marcia sells 10 dresses per month. her monthly revenue is $5,000 the marginals cost $400. in order to maximise profits Marcia should?
1
Expert's answer
2011-12-08T09:01:46-0500
To maximize profits MR = P = MC. The price P = TR/Q = 5000/10 = $500. Marcia
should reduce her price by $100, then MR = MC = P = $400 and she will maximize
her profit.